OpenStudy (dessyj1):

What are the units of the energy of a photon and the energy of light? I know light intensity is the number of photons that corresponds with the wavelength so from the information I would guess the unit would be number of waves per second.

OpenStudy (anonymous):

Energy of a photon=hv h=Planck constant. Its units are J*s or joules times seconds v=frequency. Its units are 1/s Energy of a light is energy of photon, since photon is a packet of light. Do you mean speed of light?

OpenStudy (dessyj1):

I understand now. I was never given the units to Planck's constant. I was just give a number.

OpenStudy (anonymous):

The question is just asking for units Planck constant is 6.626 x 10^-24 J*s btw

OpenStudy (dessyj1):

Yeah, i was not shown the j*s - i was only shown the number. So when i would do calculations with planck's number the units would not cancel completely because i did not know the j*s.

OpenStudy (anonymous):

You do not need numbers to figure out the units. I give you explanation, but not the answer. (I know the answer) You can figure the answer by my info. Example problem: Velocity=distance/time Units of distance=m Time=s Velocities units would be m/s Just wondering if you can figure it out.

OpenStudy (photon336):

\[E = hv \] \[E = \frac{ hc }{ \lambda }\]

OpenStudy (dessyj1):

I understand what you are saying. I did not know the units for planck's constant so when I would try a practice problem like this one, it would go like this. What is the energy of a photon of green light with a 520nm wavelength? \[E _{p}= \frac{ hc }{ \lambda } \] \[E _{p}= (6.626x10^{-34})(2.998x10^{17}nm/s)(\frac{1}{ 520nm })\] so, at the end of the day I would be left with a unit of 1/s(which did not make sense to me) because I did not know the full mangitude of planck's.

OpenStudy (anonymous):

Do you know the units for the energy of a photon now? If you do, then I can continue to solving problems with you.

OpenStudy (dessyj1):

Yes, it is in joules.

OpenStudy (dessyj1):

If I modified the question and said there were 10 photons in the wave, what would be the answer.

OpenStudy (dessyj1):

Are photons measured in mols?

OpenStudy (anonymous):

You are right. It is in joules. Modify which question?

OpenStudy (dessyj1):

What is the overall energy of green light with a 520nm wavelength that contains 10 photons?

OpenStudy (anonymous):

You would multiply it by 10 since ten more photons means then more energy. E=hv/wavelength only applies to one photon

OpenStudy (anonymous):

ten more energy*

OpenStudy (dessyj1):

would the units still be just joules?

OpenStudy (anonymous):

Yes. Lets say a photon has an energy of 500J 10(500J)=5000J

OpenStudy (dessyj1):

Okay. Thank you!

OpenStudy (dessyj1):

How would you measure photons?

OpenStudy (dessyj1):

Or is it very discrete?